# TLAC: Two-stage LMM Augmented CLIP for Zero-Shot Classification

### v1

@@ Abstract
Contrastive Language-Image Pretraining (CLIP) has shown impressive zero-shot performance on image classification. However, state-of-the-art methods often rely on fine-tuning techniques like prompt learning and adapter-based tuning to optimize CLIP’s performance. The necessity for fine-tuning significantly limits CLIP’s adaptability to novel datasets and domains. This requirement mandates substantial time and computational resources for each new dataset. To overcome this limitation, we introduce simple yet effective training-free approaches, Single-stage LMM Augmented CLIP (SLAC) and Two-stage LMM Augmented CLIP (TLAC), that leverages powerful Large Multimodal Models (LMMs), such as Gemini, for image classification. The proposed methods leverages the capabilities of pre-trained LMMs, allowing for seamless adaptation to diverse datasets and domains without the need for additional training. Our approaches involve prompting the LMM to identify objects within an image. Subsequently, the CLIP text encoder determines the image class by identifying the dataset class with the highest semantic similarity to the LLM predicted object. We evaluated our models on 11 base-to-novel datasets and they achieved superior accuracy on 9 of these, including benchmarks like ImageNet, SUN397 and Caltech101, while maintaining a strictly training-free paradigm. Our overall accuracy of 83.44% surpasses the previous state-of-the-art few-shot methods by a margin of 6.75%. Our method achieved 83.6% average accuracy across 13 datasets, a 9.7% improvement over the previous 73.9% state-of-the-art for training-free approaches. Our method improves domain generalization, with a 3.6% gain on ImageNetV2, 16.96% on ImageNet-S, and 12.59% on ImageNet-R, over prior few-shot methods.
@@ 2.2 Two-stage LMM Augmented CLIP (TLAC)
3 Experiments
3.1 Generalization from Base-to-Novel classes
3.2 Domain Generalization
3.3 Implementation Details
@@ 6 Limitations
Despite the aforementioned advantages, our method also exhibits certain limitations. For instance, it underperforms compared to fine-tuned CLIP models on the OxfordPets dataset. This performance gap can be attributed to the differing training paradigms. LMMs, trained on extensive and diverse datasets, develop more generalized object representations. Conversely, fine-tuned models learn dataset-specific correlations, providing an advantage on specialized datasets like OxfordPets. Consequently, highly specific datasets tend to yield lower accuracy with LMMs. A similar trend is observed with the EuroSAT dataset, which comprises low-resolution (64x64 pixel) aerial images categorized into 10 scene classes, such as River. The limited image resolution poses a challenge for LMMs in accurately identifying or describing the scene. Fine-tuned models, however, demonstrate superior accuracy on this dataset due to their training on class-specific features. Therefore, datasets characterized by either high specialization (e.g., OxfordPets) or low image resolution (e.g., EuroSAT) currently favor fine-tuned CLIP models over LMMs.
@@ 7.4 Large MultiModal Models (LMMs)
The large multimodal models mentioned above are closed-source, meaning their precise architecture and training data are not publicly disclosed. In contrast, open-source models make their architecture and trained weights publicly available. Pixtral-12B [2] is a 12-billion-parameter multimodal language model that excels in understanding both images and text. Qwen2-VL [41] introduces a dynamic resolution mechanism for processing images. LLaVA (Large Language and Vision Assistant) [20] is an end-to-end trained LMM that combines CLIP with the Vicuna [6]. MiniGPT-4 [51] aligns a frozen visual encoder with a frozen large language model (Vicuna) to replicate GPT-4’s advanced multimodal abilities. Llama [9] is a new family of language models. While Llama 3 [37] and 3.1 [9] were language models, Llama 3.2 introduced multimodal capabilities.

### v2

@@ Abstract
Contrastive Language-Image Pretraining (CLIP) has shown impressive zero-shot performance on image classification. However, state-of-the-art methods often rely on fine-tuning techniques like prompt learning and adapter-based tuning to optimize CLIP’s performance. The necessity for fine-tuning significantly limits CLIP’s adaptability to novel datasets and domains. This requirement mandates substantial time and computational resources for each new dataset. To overcome this limitation, we introduce simple yet effective training-free approaches, Single-stage LMM Augmented CLIP (SLAC) and Two-stage LMM Augmented CLIP (TLAC), that leverages powerful Large Multimodal Models (LMMs), such as Gemini, for image classification. The proposed methods leverages the capabilities of pre-trained LMMs, allowing for seamless adaptation to diverse datasets and domains without the need for additional training. Our approaches involve prompting the LMM to identify objects within an image. Subsequently, the CLIP text encoder determines the image class by identifying the dataset class with the highest semantic similarity to the LLM predicted object. Our models achieved superior accuracy on 9 of 11 base-to-novel datasets, including ImageNet, SUN397, and Caltech101, while maintaining a strictly training-free paradigm. Our TLAC model achieved an overall accuracy of 83.44%, surpassing the previous state-of-the-art few-shot methods by a margin of 6.75%. Compared to other training-free approaches, our TLAC method achieved 83.6% average accuracy across 13 datasets, a 9.7% improvement over the previous methods. Our Code is available at https://github.com/ans92/TLAC
@@ 2.2 Two-stage LMM Augmented CLIP (TLAC)
3 Experiments
3.1 Generalization from Base-to-Novel classes
3.2 Domain Generalization
3.3 Implementation Details
@@ 6 Limitations
While the aforementioned advantages are notable, our method also presents certain limitations. A primary concern pertains to the LMM model, specifically whether the observed accuracy is attributable to our unique model configuration or the inherent power of the Gemini architecture. Furthermore, the potential for data contamination exists, raising concerns that the LMM model may have been exposed to the datasets employed for evaluation. To mitigate the first limitation, future studies could investigate the performance of different LMM models to isolate the contribution of our specific configuration from that of the underlying backbone. Regarding the second limitation, the development of a novel benchmark dataset, previously unseen by the LMM, would be a valuable contribution. We recommend that future research endeavors focus on these areas.
Another limitation is that LMMs tend to show lower accuracy on highly specific datasets. For example, our model underperforms fine-tuned CLIP on OxfordPets because fine-tuning allows models to learn dataset-specific correlations, a significant advantage on specialized datasets like OxfordPets. Similarly, we observe this trend with EuroSAT, a low-resolution (64x64 pixel) aerial image dataset with 10 scene classes (e.g., River). Consequently, datasets with high specialization (e.g., OxfordPets) or low image resolution (e.g., EuroSAT) currently favor fine-tuned CLIP models over LMMs.
@@ 7.4 Large MultiModal Models (LMMs)
The large multimodal models mentioned above are closed-source, meaning their precise architecture and training data are not publicly disclosed. In contrast, open-source models make their architecture and trained weights publicly available. Pixtral-12B [2] is a 12-billion-parameter multimodal language model. Qwen2-VL [41] introduces a dynamic resolution mechanism for processing images. LLaVA (Large Language and Vision Assistant) [20] is an end-to-end trained LMM that combines CLIP with the Vicuna [6]. MiniGPT-4 [51] aligns a frozen visual encoder with a frozen large language model (Vicuna) to replicate GPT-4’s advanced multimodal abilities. Llama [9] is a new family of language models. While Llama 3 [37] and 3.1 [9] were language models, Llama 3.2 introduced multimodal capabilities.